# Nonlinear skin modes and fixed-points

### v1

@@ II Model
The SIBC spectra for both the linear and nonlinear systems fill the same disk, bounded by the linear PBC loop in the complex energy plane, but the corresponding OBC spectra do not share this similarity. In the linear system, the OBC spectrum is confined to the interval of $\mp 2\sqrt{\gamma}$ on the real energy axis, with its width ($4\sqrt{\gamma}$) being independent of the system size, implying that it is a subset of the corresponding SIBC spectrum. However, in the nonlinear case, OBC spectrum is intriguingly not bounded by the SIBC spectrum and its width decreases with increasing lattice size, eventually approaching that of the SIBC on the real energy axis. In other words, OBC skin modes exist with energy $E$ outside the range defined by ${\mp}E_{c}$. The key factor behind this intriguing behavior is that, for either $E>E_{c}$ or $E<-E_{c}$, $\psi_{n}$ initially decays and becomes zero at a certain point $n_{0}$ (satisfying OBC for the system with $N=n_{0}-1$) but then slowly increases to a nonzero fixed point and remains at that fixed point, violating SIBC. Fig. 2(a) confirms this by presenting $\psi_{N+1}$ as a function of $\psi_{1}$ at $E=1.3>E_{c}$ for three different system sizes. We see that OBCs are satisfied at specific values of $\psi_{1}$, which are almost the same for the three lattice sizes (indistinguishable by the naked eye in the figure). Fig. 2(b) shows the density plots for some values of $\psi_{1}$, revealing that OBC is satisfied at $N=3$, whereas SIBC is not, as the field amplitude reaches a fixed point at large values of $n$. To this end, we would like to say that $N_{c}$ does not take finite values when $E$ is complex, suggesting that the nonlinear OBC spectrum is real valued, similar to the linear OBC spectrum.
@@
Figure 2: (a) The condition $\psi_{N+1}=0$ is satisfied at specific points that are very close to each other for different lattice sizes: $N=5$ (blue), $N=10$ (red), and $N=100$ (black). (b) $\psi_{n}$ for 3 different $\psi_{1}$ up to $n=60$ show that the field amplitude takes zero value (at $n=4$) before reaching a nonzero fixed point. This means that OBC modes are available at $N=3$, but the SIBCs are not at this energy. The common parameters are $E=1.3>E_{c}$, $\gamma=0.2$, and $g_{1}=1$ (only Kerr-type nonlinearity). For the insets, the nonlinear strengths are $\alpha=\beta=g_{1}=20g_{2}=1$, $g_{3}=0$
@@
Figure 3: (a) $|\psi_{n}|$ for the localized OBC modes in the presence of a single coupling impurity at the edge when $W_{1}=1$, $\psi_{1}=10^{-14}$, $E=2$ and $N=200$ ($N=100$ in the inset). (b) The dark and anti-dark soliton under PBC in the system with $\psi_{1}=0.447$, $1503p=p^{\prime}=$ 150, $W_{2}=-W_{1}=0.8$ and $E=1.4$. The width of the dark soliton can be increased by placing the second impurity next to the first one (see the inset at $p=p^{\prime}-1=50$, and $W_{1}=1$ and $W_{2}=-0.2002$). The common parameters are $\gamma=0.2$ and $g_{1}=1$.
@@ III Impurity induced modes
Let us introduce a coupling impurity with strength $W_{1}$ at site $p$ such that $J_{R}$ in Eq. (1) is modified to ${J_{R}\rightarrow\gamma+W_{1}\delta_{n,p}}$ ($J_{L}=1$). Consider $E>E_{c}$, where $\psi_{n}$ converges to a nonzero fixed value. When the impurity with an appropriate strength is introduced at the right edge, it cause the field amplitude to drop to zero value from the nozero fixed point at that edge. Conversely, when the impurity is located in the bulk, it perturbs the field amplitude around itself, while the field amplitude away from the impurity in both directions is equal to the nonzero fixed point. The former one may lead to localized modes, while the latter one to dark solitons.
@@
We start with the case where the impurity is located at the right edge. Since the initial value $\psi_{1}$ does not change the nonzero fixed point, we propose assuming a very small value for $\psi_{1}$, such as $\psi_{1}{<}10^{-10}$. In this case, the field amplitude is practically zero near the left edge, and reaches the nonzero fixed value away from the left edge, and remains constant towards the right edge, thus allowing us to construct localized modes at the right edge. Note that the OBC is satisfied with an appropriate choice of the impurity strength $W_{1}$. As the nonzero fixed value varies with $E$, a family of localized modes is generated until $E$ becomes sufficiently large for the system to enter into a chaotic regime. As an illustration, Fig. 3(a) plots $|\psi_{n}|$ when $N=200$ and $\psi_{1}{=}10^{-14}$. Notably, the specific lattice site where the field amplitude reaches the nonzero fixed value remains the same regardless of the system size as can be seen in the inset in Fig. 3(a) for $N=100$. Therefore, these modes are neither skin nor scale free localized modes whose localization lengths remain constant or change proportionally with the system size, respectively.
@@
A single coupling impurity may also lead to dark or anti-dark soliton solutions under PBC. A dark (anti dark) soliton is a localized drop (increase) in the density on the uniform background. To obtain them, we place the impurity in the bulk and perform the iteration with $\psi_{1}=a_{0}$ (period one orbit) and the corresponding energy such that the extended PBC state with uniform density would be obtained in the absence of the impurity. Therefore, the field amplitude at the left side of the impurity is not changed: $\psi_{n}=a_{0}$ when $n<p$. However, the field amplitude will either increase or dip at the impurity location, depending on the sign of $W_{1}$, and then returns to the fixed point after a few lattice sites, remaining constant towards the right edge. In this way, we obtain dark solitons when $W_{1}>0$ and anti-dark solitons when $W_{1}<0$. Additionally, if the system contains two well-separated impurities at sites $p$ and $p^{\prime}$ with opposite strengths, $W_{2}=-W_{1}$, we can generate a pair of dark and anti-dark solitons, as illustrated in Fig. 3 (b). Introducing additional well-separated impurities into the system allows for the formation of multiple dark and anti-dark solitons. On the other hand, placing the two impurities on neighboring lattice points, $p^{\prime}=p\mp 1$, allows us to adjust both the dip point and the width of the dark soliton. As shown in the inset in Fig. 3 (b), the density drops to zero at the impurity position and then returns to its fixed point after a certain distance, leading to a wider dark soliton. Adding more neighboring impurities allows us to obtain much wider dark solitons.

### v2

@@ II Model
The SIBC spectra for both the linear and nonlinear systems fill the same disk, bounded by the linear PBC loop in the complex energy plane, but the corresponding OBC spectra do not share this similarity. For the linear system, the OBC spectrum is confined to the real energy interval $\mp 2\sqrt{\gamma}$, indicating that it is a subset of the corresponding SIBC spectrum. However, for the nonlinear case, OBC spectrum is not bounded by the SIBC spectrum and its width is size-dependent. In other words, nonlinear OBC skin modes exist at real energy values $E$ outside the range of ${\mp}E_{c}$. The key factor behind this intriguing behavior is that, for ${E>E_{c}}$, $\psi_{n}$ may initially decay and become zero at a certain point (satisfying OBC) before eventually rising to the nonzero fixed point (breaking SIBC). In other words, if ${\psi_{N+1}=0}$, then ${\psi_{\infty}=0}$ for the linear system. However, in the nonlinear system, it is possible for ${\psi_{\infty}}$ to be a nonzero fixed point, even when ${\psi_{N+1}=0}$. Fig. 2 illustrates this for the system with only Kerr-type nonlinearity, while the insets correspond to the system with different values of the nonlinear parameters. Fig. 2(a) shows $\psi_{N+1}$ as a function of $\psi_{1}$ at ${E=1.3>E_{c}}$ for three different system sizes. We see that the condition ${\psi_{N+1}=0}$ is satisfied at specific values of $\psi_{1}$, which are almost the same for the three lattice sizes (indistinguishable by the naked eye in the figure). Fig. 2(b) shows the density plots for three specific modes with the same energy but different $\psi_{1}$. These modes are OBC modes for $N=3$ (${\psi_{4}=0}$) but not SIBC modes, as $\psi_{n}$ eventually reaches the same fixed point.
@@
Figure 2: (a) The condition $\psi_{N+1}=0$ is satisfied at specific points that are very close to each other for different lattice sizes: $N=5$ (blue), $N=10$ (red), and $N=100$ (black). (b) $|\psi_{n}|$ for three different values of $\psi_{1}$ up to $n=60$, where ${\psi_{n}}$ becomes zero at ${n=4}$. The common parameters are ${E=1.3>E_{c}}$, ${\gamma=0.2}$, and $g_{1}=1$ (only Kerr-type nonlinearity). For the insets, the nonlinear strengths are ${\alpha=\beta=g_{1}=20g_{2}=1}$ and ${g_{3}=0}$.
@@
Figure 3: (a) $|\psi_{n}|$ in the presence of a single coupling impurity at the edge when $W_{1}=1$, ${\psi_{1}=10^{-14}}$, $E=2$ and $N=200$ ($N=100$ in the inset). (b) The dark and anti-dark soliton at ${\psi_{0}=\psi_{1}=0.461}$, $1503p=p^{\prime}=$ 150, $W_{2}=-W_{1}=0.8$ and $E=1.4$. The width of the dark soliton can be increased by placing the second impurity next to the first one (see the inset at $p=p^{\prime}-1=50$, and $W_{1}=1$ and $W_{2}=-0.2002$). The common parameters are $\gamma=0.2$, $g_{1}=2g_{2}=5g_{3}=\beta=1$ and ${\alpha=0}$.
@@ III Impurity induced modes
Consider ${E>E_{c}}$, where $\psi_{n}$ converges to a nonzero fixed value. Next, we introduce a coupling impurity with strength $W_{1}$ at site ${p>N_{c}}$ such that $J_{R}$ in Eq.(II) is modified to ${J_{R}\rightarrow\gamma+W_{1}\delta_{n,p}}$ (${J_{L}=1}$). When the impurity with an appropriate strength is introduced at the right edge, it causes the field amplitude to drop to zero value from the nonzero fixed value at that edge. Conversely, when the impurity is located far from the right edge, it perturbs the field amplitude only locally without changing the nonzero fixed value towards the right edge. The former one may lead to localized modes under OBC, while the latter one to dark solitons under PBC, as shown below.
@@
We start with the case where the impurity is located at the right edge for the lattice under OBC. Since the initial value $\psi_{1}$ does not change the nonzero fixed point, we propose assuming a very small value for $\psi_{1}$, such as ${\psi_{1}{<}10^{-10}}$ (${\psi_{0}=0}$). In this case, the field amplitude remains practically zero near the left edge until it grows sharply to a nonzero fixed value at around the point ${N_{c}}$, and then it remains constant towards the right edge. Note that the OBC at the right edge ${\psi_{N+1}=0}$ is satisfied with an appropriate choice of the impurity strength $W_{1}$. In this way, we construct novel localized modes at the right edge. As the nonzero fixed point varies with $E$, a family of such localized modes is generated by varying $E$ until $E$ becomes sufficiently large for the system to enter into a chaotic regime. The point ${N_{c}}$ does not change with the system size, hence these modes extend nearly from the point ${N_{c}}$ to the right edge. Therefore, these modes are different from the left-localized skin modes as they are localized at the right edge and their extension changes with system size. They are also different from scale-free modes localized at the right edge, which are typically given by ${\psi_{n}\approx e^{cx/N}}$ in the linear system, where $c$ is a constant. As an illustration, Fig. 3(a) plots $|\psi_{n}|$ at ${E=2}$, ${\psi_{1}{=}10^{-14}}$, ${W_{1}=1}$ when the lattice with ${N=200}$ has a saturable cubic-quintic-sextic nonlinearity. The specific lattice site where the field amplitude reaches the nonzero fixed value remains the same regardless of the system size as can be seen in the inset in Fig. 3(a) for $N=100$. As can be seen, these modes are neither skin nor scale free localized modes whose localization lengths remain constant or change proportionally with the system size, respectively.
@@
A single coupling impurity may also lead to dark or anti-dark soliton solutions under PBC. A dark (anti dark) soliton is a localized drop (increase) in the density on the uniform background. To obtain them, we place the impurity far from the edges and perform the iteration with ${\psi_{1}=a_{0}}$ (period one orbit) and ${E>E_{c}}$ such that the extended PBC state with uniform density would be obtained in the absence of the impurity. Therefore, the field amplitude at the left side of the impurity is not changed: ${\psi_{n}=a_{0}}$ when ${n{\leq}p}$. However, the field amplitude will either increase or dip at ${n=p+1}$, depending on the sign of $W_{1}$: ${\psi_{p+1}=(1-W_{1})a_{0}}$. Recall that the fixed point is stable against small fluctuations. Therefore the field amplitude returns to the fixed point after a few lattice sites, remaining constant towards the right edge. In this way, we obtain dark solitons when $W_{1}>0$ and anti-dark solitons when ${W_{1}<0}$. Additionally, if the system contains two well-separated impurities at sites $p$ and $p^{\prime}$ with opposite strengths, ${W_{2}=-W_{1}}$, we can generate a pair of dark and anti-dark solitons, as illustrated in Fig. 3 (b), where the impurities are located at ${p^{\prime}=3p=150}$. Introducing additional well-separated impurities into the system allows for the formation of multiple dark and anti-dark solitons. On the other hand, placing the two impurities on neighboring lattice points, $p^{\prime}=p\mp 1$, allows us to adjust both the dip point and the width of the dark soliton. In the inset in Fig. 3 (b), we consider the two adjacent impurities with appropriate strengths. As seen, the density drops to zero at the impurity position and then returns to its fixed point after a certain distance, leading to a wider dark soliton. Adding more neighboring impurities allows us to obtain much wider dark solitons. It is worth noting that skin modes also exists when ${E{\leq}E_{c}}$, as the zero fixed point is stable against the coupling impurity. This indicates that skin modes and dark solitons coexist within the system, depending on the energy.